# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v4

@@ Abstract
Implicit Neural Representation (INR) has been successfully employed for Arbitrary-scale Super-Resolution (ASR). However, INR-based models need to query the multi-layer perceptron module numerous times and render a pixel in each query, resulting in insufficient representation capability and computational efficiency. Recently, Gaussian Splatting (GS) has shown its advantages over INR in both visual quality and rendering speed in 3D tasks, which motivates us to explore whether GS can be employed for the ASR task. However, directly applying GS to ASR is exceptionally challenging because the original GS is an optimization-based method through overfitting each single scene, while in ASR we aim to learn a single model that can generalize to different images and scaling factors. We overcome these challenges by developing two novel techniques. Firstly, to generalize GS for ASR, we elaborately design an architecture to predict the corresponding image-conditioned Gaussians of the input low-resolution image in a feed-forward manner. Each Gaussian can fit the shape and direction of an area of complex textures, showing powerful representation capability. Secondly, we implement an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to render super-resolved images by sampling discrete RGB values from the predicted continuous Gaussians. Via end-to-end training, our optimized network, namely GSASR, can perform ASR for any image and unseen scaling factors. Extensive experiments validate the effectiveness of our proposed method. The code and models are available at https://github.com/ChrisDud0257/GSASR.
@@ 1 Introduction
Image Super-Resolution (SR) has been a challenging inverse problem in computer vision for many years, and a variety of deep learning-based (SR) methods have been developed in the past decade [46, 58, 31, 5, 43, 47, 56, 50, 51, 55, 6]. However, most of the existing SR models [63, 64, 45, 30, 61, 9, 15, 29, 59] are designed to operate only under integer or fixed scaling factors, i.e., ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4, due to the fact that they perform resolution upsampling by processing discrete pixels with de-convolutional layers [15] or pixel-shuffle layers [40]. As a result, it is hard for them to perform SR with decimal or even arbitrary scaling factors. Although the existing SR models could achieve arbitrary-scale zooming via adopting bicubic or bilinear interpolation operations, they would inevitably lead to unsatisfied or over-smoothing super-resolved results.
Inspired by the development of Implicit Neural Representation (INR) [37], Chen et al. presented LIIF [10] to address the problem of Arbitrary-scale SR (ASR) by making use of the continuity of Multi-Layer Perceptron (MLP) to query the RGB values of a high-resolution (HR) image and feeding its index to the MLP network. The following works try to improve the ASR performance with local feature estimators [27], normalization flow [54] or scale-aware attention modules [2]. Most of the existing INR-based ASR models [10, 27, 54, 2] rely heavily on MLP layers to predict RGB pixels. However, this implicit function lacks enough representation capacity, as it renders the RGB color of a single pixel in each query. INR-based models need to query the MLP layers numerous times, resulting in very low computational efficiency, especially for those models [2] who scale up the MLP layers and employ complex attention modules.
@@
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [22] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution). Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [32, 63]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embeddings to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [34] so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the continuous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
The advantages of our proposed GSASR are three-folds. (1) Since each Gaussian contributes to a large area, GSASR outputs region-level predictions, showing much more powerful representation capability than INR-based models, which only render a single pixel during each query. (2) The attributes of Gaussians are explicitly predicted to ensure better reconstruction results. Specifically, the predicted position enables Gaussians to concentrate on areas with complex textures, while the predicted standard deviation and correlation coefficient indicate the shape and direction of Gaussians, allowing Gaussian to fit the shape of complex textures. (3) Benefited from the developed efficient scale-aware 2D rasterization technique, GSASR is much faster than the state-of-the-art INR-based models. As shown in Fig. 1 and Tab. 2, under the scaling factor ×$\times 12$, GSASR could super-resolve an image to $720\times 720$ resolution within only $91$ ms, while the state-of-the-art INR-based model such as CiaoSR [2] will cost nearly $540$ ms.
@@ 2 Related Work
Arbitrary-scale SR Methods. The interpolation operators can be used to continuously predict neighborhood pixels at any magnification factors, yet they will sacrifice the image quality and obtain blurry results. Meta-SR [20] designsing a meta-upscaling module that leverages DNN to predict HR details at arbitrary-scales from LR neighborhood embeddings. Wang et al. [44] performed arbitrary upsampling by introducing dynamic filters generated based on the scaling factor. Inspired by the success of INR-based methods [37] in 3D tasks, LIIF [10] introduces INR to complete the ASR task by learning an implicit mapping function between a pixel’s RGB value and its index. ITSRN [52] makes use of a transformer block and implicit position embedding to predict the super-resolved images. LTE [27] employs an implicit function to describe image features in Fourier domain. SRNO [49] adopts a kernel integral operator to fit the implicit function. LINF [54] estimates a local implicit function with normalizing flow to model the naturalness of textures. CLIT [7] employs a cascaded cross-scale attention block with frequency encoding layers to sense neighborhood features. CiaoSR [2] introduces scale-aware attention-in-attention modules to extend the receptive field to promote the ASR performance. LMF [18] introduces a latent modulated function to compress the feature and accelerate the rendering process at the price of representation accuracy.
Very recently, GaussianSR [19] was proposed to employ GS for the ASR task. However, it only utilizes the pixel positions of LR image to substitute the center positions of Gaussians, heavily limiting the representation capability of Gaussians. In contrast, our proposed GSASR predicts all of the parameters of Gaussians, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
@@ 3 Method
2D GS Representation. Our key insight is to represent an image using continuous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R},\mu_{y}~{}\in\mathbb{R}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty),\sigma_{y}~{}\in(0,\infty)\}$, correlation coefficient $\rho~{}\in[-1,1]$, and normalized peak RGB value $c~{}\in[0,1]$. We define the RGB color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
$G(x,y)=\alpha\cdot c\cdot f(x,y).$ (1)
@@
$\vspace{-0.8em}\begin{split}f(x,y)=&{(2\pi\sigma_{x}\sigma_{y}\sqrt{1-\rho^{2}})^{-1}}exp[{-\frac{1}{2(1-\rho^{2})}}\times\\ &{(\frac{\Delta x^{2}}{\sigma_{x}^{2}}-\frac{2\rho\Delta x\Delta y}{\sigma_{x}\sigma_{y}}+\frac{\Delta y^{2}}{\sigma_{y}^{2}})}]\end{split}$ (2)
@@
$I_{LR}(x,y)=\sum_{i=1}^{N}G_{i}(x,y).\vspace{-0.6em}$ (3)
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).\vspace{-0.8em}$ (4)
Fast Rasterization. In Eq. 4, simply rendering an SR image by querying each pixel from all 2D Gaussians leads to a complexity of $\mathcal{O}(s^{2}HWN)$, which is too high for high-resolution images. Actually, a Gaussian generally focuses on a local area and its contribution to pixel values decays rapidly with the increase of distance. Therefore, we introduce a rasterization ratio $r\leq 1$ to control the rendering range of each Gaussian. Specifically, we handle all Gaussians in parallel and only render the pixels that are close enough to the Gaussian centers, greatly reducing the computational complexity to $\mathcal{O}(r^{2}s^{2}HWN)$ and making our algorithm efficient. To achieve a better balance between performance and efficiency, we set the ratio as $r=0.1$, which could significantly speed up the rasterization process and avoid missing major Gaussian responses. Note that our rasterization process is differential, which can be seamlessly integrated into neural network training for end-to-end optimization. Algorithm 1 describes the detailed steps of our scale-aware rasterization process. This algorithm is implemented via CUDA C++, which is GPU-friendly and achieves fast speed and low memory requirements.
@@ 6       For each pixel $(x,y)$ in $I_{SR}$ do
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [32] as the feature extraction encoder and are tested on DIV2K and LSDIR [28] datasets [41] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. PSNR/SSIM metrics are computed on the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Conventional 3D GS methods [22] directly optimize 3D Gaussians to fit a specific scene. As a result, they lack the generalization capability for the ASR tasks, where the model should be generalized to unseen images. To overcome this limitation, we introduce a novel network architecture, which conditions 2D Gaussians on LR input images, generating content-aware Gaussian representations with enhanced generalization capabilities. Our proposed method, namely GSASR, is illustrated in Figure. 2. It begins with an off-the-shelf encoder (such as EDSR-Baseline [32], RDN [64]) that extracts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ from the LR input $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$. The deep features $\mathbf{F}$ are input to the subsequent Gaussian generation process. The key of GSASR lies in the proposed Gaussian decoder, which receives image feature $\mathbf{F}$ and learnable Gaussian embeddings as inputs, and outputs 2D Gaussians $\mathcal{G}$ to depict the input image in continuous Gaussian space. Our Gaussian decoder consists of a condition injection block, a Gaussian interaction block, and a Gaussian primary head. Finally, the Gaussian decoder output is rasterized into an SR image.
@@
Gaussian Embedding. In Gaussian decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. As for the number of Gaussian embeddings $N$, a straightforward idea is to set it to a fixed number so that the $N$ independent learnable Gaussian embeddings interact with image features to output Gaussians $\mathcal{G}$. However, this approach will either be ineffective or lead to low computational efficiency. As the LR images exhibit variable sizes in the test phase, rendering a fixed number of Gaussians will either waste computational resources when handling LR image with very small size, or lead to deficient representation when processing larger LR images. Another possible strategy is to set $N\propto s$ under different scaling factord $s$. However, it requires the Gaussian decoder to handle varying sizes of Gaussian embeddings, while the size of LR features is fixed in the training stage. In addition, as the scaling factor increases, more Gaussians are introduced, leading to longer inference time and higher GPU memory cost.
@@
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,\vspace{-0.5em}$ (5)
@@
Figure 4: Visualization of GSASR and the competing methods under ×$\times 4$ scaling factor with EDSR [32] feature extraction backbone. The competing methods result in blurry details, while GSASR generates much clear textures.
@@
Figure 5: Visualization of GSASR and the competing methods under ×$\times 6$ scaling factor with RDN [64] feature extraction backbone. The competing methods produce blurry textures, while GSASR generates much clearer contents.
Gaussian Primary Head. We employ five disentangled heads to convert Gaussian embeddings (output from Gaussian Interaction Block) $\mathbf{E}$ into the corresponding five properties of 2D Gaussians $\mathcal{G}$. To be more specific, the $i$-th Gaussian embedding $\mathbf{E}_{i}$ is fed into five lightweight MLP blocks, each of which contains three linear layers and two ReLU activation functions, to output its opacity $\alpha_{i}$, color $c_{i}$, stand deviation $\sigma_{i}$, reference offset $o_{i}$, and correlation $\rho_{i}$. Note that the position is obtained by adding the reference position and predicted offset $\mu_{i}=p_{i}+o_{i}$. The five property heads are decoupled so that different heads can be adapted to predict specific properties. We apply the tanh function to ensure $\rho\in[-1,1]$, and use the sigmoid function for $\{\alpha,c,\sigma\}$ to ensure that their values are physically meaningful. No activation function is applied to $o$ since the positions of Gaussians in 2D images are expected to be free. Finally, the 2D Gaussians $\mathcal{G}$ are calculated from Eq. 1, Eq. 2 with the predicted five properties $\{\alpha,\mu,c,\sigma,\rho\}$.
@@
Scale Computational Cost and Performance Backbone: EDSR-baseline

### v5

@@ Abstract
Implicit Neural Representations (INR) have been successfully employed for Arbitrary-scale Super-Resolution (ASR). However, INR-based models need to query the multi-layer perceptron module numerous times and render a pixel in each query, resulting in insufficient representation capability and low computational efficiency. Recently, Gaussian Splatting (GS) has shown its advantages over INR in both visual quality and rendering speed in 3D tasks, which motivates us to explore whether GS can be employed for the ASR task. However, directly applying GS to ASR is exceptionally challenging because the original GS is an optimization-based method through overfitting each single scene, while in ASR we aim to learn a single model that can generalize to different images and scaling factors. We overcome these challenges by developing two novel techniques. Firstly, to generalize GS for ASR, we elaborately design an architecture to predict the corresponding image-conditioned Gaussians of the input low-resolution image in a feed-forward manner. Each Gaussian can fit the shape and direction of an area of complex textures, showing powerful representation capability. Secondly, we implement an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to render super-resolved images by sampling discrete RGB values from the predicted continuous Gaussians. Via end-to-end training, our optimized network, namely GSASR, can perform ASR for any image and unseen scaling factors. Extensive experiments validate the effectiveness of our proposed method. The code and models are available at https://github.com/ChrisDud0257/GSASR.
@@ 1 Introduction
Image Super-Resolution (SR) has been a challenging inverse problem in computer vision for many years, and a variety of deep learning-based (SR) methods have been developed in the past decade [47, 60, 32, 5, 44, 48, 58, 51, 52, 57, 6, 7, 56]. However, most of the existing SR models [65, 66, 46, 31, 63, 10, 16, 30, 61] are designed to operate only under integer or fixed scaling factors, i.e., ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4, due to the fact that they perform resolution upsampling by processing discrete pixels with de-convolutional layers [16] or pixel-shuffle layers [41]. As a result, it is hard for them to perform SR with decimal or even arbitrary scaling factors. Although the existing SR models could achieve arbitrary-scale zooming via adopting bicubic or bilinear interpolation operations, they would inevitably lead to unsatisfied or over-smoothing super-resolved results.
Inspired by the development of Implicit Neural Representations (INR) [38], Chen et al. presented LIIF [11] to address the problem of Arbitrary-scale SR (ASR) by making use of the continuity of Multi-Layer Perceptron (MLP) to query the RGB values of a high-resolution (HR) image and feeding its index to the MLP network. The following works try to improve the ASR performance with local feature estimators [28], normalization flow [55] or scale-aware attention modules [2]. Most of the existing INR-based ASR models [11, 28, 55, 2] rely heavily on MLP layers to predict RGB pixels. However, this implicit function lacks enough representation capacity, as it renders the RGB color of a single pixel in each query. INR-based models need to query the MLP layers numerous times, resulting in very low computational efficiency, especially for those models [2] who scale up the MLP layers and employ complex attention modules.
@@
Instead of overfitting the model to a specified LR image, we switch from the optimization-based manner of original 3D GS [23] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution). Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [33, 65]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embeddings to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [35] so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the continuous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as conditions. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
The advantages of our proposed GSASR are threefold. (1) Different from previous INR-based models, which query each HR pixel exclusively and lack interaction among pixels, GSASR employs 2D Gaussians to perform region-level representation, providing a much more powerful representation capability. (2) In GSASR, the Gaussians’ centers are learnable and input adaptive so that the Gaussians can concentrate on areas with complex textures. In addition, the standard deviation and correlation coefficient of Gaussians indicate their shape and direction, allowing Gaussians to fit the different shapes of different objects and scenes. (3) Taking advantage of our developed efficient scale-aware 2D rasterization techniques, GSASR is much faster than state-of-the-art INR-based models. As shown in Fig. 1 and Tab. 2, under the scaling factor ×12$\times$ 12× 12, GSASR could super-resolve an image to $\times 720$ 720 resolution within only 919191 ms, while the state-of-the-art INR-based model such as CiaoSR [2] will cost nearly 540540540 ms.
@@ 2 Related Work
Arbitrary-scale SR Methods. The interpolation operators can be used to continuously predict neighborhood pixels at any magnification factors, yet they will sacrifice the image quality and obtain blurry results. Meta-SR [21] designs a meta-upscaling module that leverages DNN to predict HR details at arbitrary-scales from LR neighborhood embeddings. Wang et al. [45] performed arbitrary upsampling by introducing dynamic filters generated based on the scaling factor. Inspired by the success of INR-based methods [38] in 3D tasks, LIIF [11] introduces INR to complete the ASR task by learning an implicit mapping function between a pixel’s RGB value and its index. ITSRN [53] makes use of a transformer block and implicit position embedding to predict the super-resolved images. LTE [28] employs an implicit function to describe image features in Fourier domain. SRNO [50] adopts a kernel integral operator to fit the implicit function. LINF [55] estimates a local implicit function with normalizing flow to model the naturalness of textures. CLIT [8] employs a cascaded cross-scale attention block with frequency encoding layers to sense neighborhood features. CiaoSR [2] introduces scale-aware attention-in-attention modules to extend the receptive field to promote the ASR performance. LMF [19] introduces a latent modulated function to compress the feature and accelerate the rendering process at the price of representation accuracy.
While the recent work GaussianSR [20] shares a similar motivation to introduce GS for the ASR task, our work has two key differences from it. First, GaussianSR places the centers of Gaussians at the pixel positions of the LR image, heavily limiting the representation capability of Gaussians. In contrast, GSASR predicts all Gaussian attributes in a free and implicit manner to unleash its expressive potential. Second, GaussianSR is slow due to its inefficient use of GS. Instead, GSASR employs an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to achieve ASR effectively and efficiently.
@@ 3 Method
[…]
$G(x,y)=\alpha\cdot c\cdot f(x,y),$ (1)
@@
$\vskip-8.00003pt\begin{split}f(x,y)=&{(2\pi\sigma_{x}\sigma_{y}\sqrt{1-\rho^{2}})^{-1}}exp[{-\frac{1}{2(1-\rho^{2})}}\times\\ &{(\frac{\Delta x^{2}}{\sigma_{x}^{2}}-\frac{2\rho\Delta x\Delta y}{\sigma_{x}\sigma_{y}}+\frac{\Delta y^{2}}{\sigma_{y}^{2}})}],\end{split}$ (2)
@@
$I_{LR}(x,y)=\sum_{i=1}^{N}G_{i}(x,y),\vskip-6.00006pt$ (3)
@@ 6 For each pixel $(x,y)$ in $I_{SR}$ do
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}),\vskip-8.00003pt$ (4)
Fast Rasterization. In Eq. 4, simply rendering an SR image by querying each pixel from all 2D Gaussians leads to a complexity of $\mathcal{O}(s^{2}HWN)$, which is too high for high-resolution images. Actually, a Gaussian generally focuses on a local area and its contribution to pixel values decays rapidly with the increase of distance. Therefore, we introduce a rasterization ratio $r\leq 1$ to control the rendering range of each Gaussian. Specifically, we handle all Gaussians in parallel and only render the pixels that are close enough to the Gaussian centers, greatly reducing the computational complexity to $\mathcal{O}(r^{2}s^{2}HWN)$ and making our algorithm efficient. To achieve a better balance between performance and efficiency, we set the ratio as $r=0.1$, which could significantly speed up the rasterization process and avoid missing major Gaussian responses. Note that our rasterization process is differentiable, which can be seamlessly integrated into neural network training for end-to-end optimization. Algorithm 1 describes the detailed steps of our scale-aware rasterization process. This algorithm is implemented via CUDA C++, which is GPU-friendly and achieves fast speed and low memory requirements.
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [33] as the feature extraction encoder and are tested on DIV2K and LSDIR [29] datasets [42] with scaling factors ×2$\times$ 2× 2, ×3$\times$ 3× 3, ×4$\times$ 4× 4, ×6$\times$ 6× 6, ×8$\times$ 8× 8, ×12$\times$ 12× 12, ×16$\times$ 16× 16, ×18$\times$ 18× 18, ×24$\times$ 24× 24, ×30$\times$ 30× 30. The best results are highlighted in red. PSNR/SSIM metrics are computed on the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Conventional 3D GS methods [23] directly optimize 3D Gaussians to fit a specific scene. As a result, they lack the generalization capability for the ASR tasks, where the model should be generalized to unseen images. To overcome this limitation, we introduce a novel network architecture, which conditions 2D Gaussians on LR input images, generating content-aware Gaussian representations with enhanced generalization capabilities. Our proposed method, namely GSASR, is illustrated in Figure. 2. It begins with an off-the-shelf encoder (such as EDSR-Baseline [33], RDN [66]) that extracts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ from the LR input $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$. The deep features $\mathbf{F}$ are input to the subsequent Gaussian generation process. The key of GSASR lies in the proposed Gaussian decoder, which receives image features $\mathbf{F}$ and learnable Gaussian embeddings as inputs, and outputs 2D Gaussians $\mathcal{G}$ to depict the input image in continuous Gaussian space. Our Gaussian decoder consists of a condition injection block, a Gaussian interaction block, and a Gaussian primary head. Finally, the Gaussian decoder output is rasterized into an SR image.
@@
Gaussian Embedding. In Gaussian decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. As for the number of Gaussian embeddings $N$, a straightforward idea is to set it to a fixed number so that the $N$ independent learnable Gaussian embeddings interact with image features to output Gaussians $\mathcal{G}$. However, this approach will either be ineffective or lead to low computational efficiency. As the LR images exhibit variable sizes in the test phase, rendering a fixed number of Gaussians will either waste computational resources when handling LR image with very small size, or lead to deficient representation when processing larger LR images. Another possible strategy is to set $N\propto s$ under different scaling factor $s$. However, it requires the Gaussian decoder to handle varying sizes of Gaussian embeddings, while the size of LR features is fixed in the training stage. In addition, as the scaling factor increases, more Gaussians are introduced, leading to longer inference time and higher GPU memory cost.
@@
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,\vskip-5.0pt$ (5)
@@
Figure 4: Visualization of GSASR and the competing methods under ×4$\times$ 4× 4 scaling factor with EDSR [33] feature extraction backbone. The competing methods result in blurry details, while GSASR generates much clear textures.
@@
Figure 5: Visualization of GSASR and the competing methods under ×6$\times$ 6× 6 scaling factor with RDN [66] feature extraction backbone. The competing methods produce blurry textures, while GSASR generates much clearer contents.
Gaussian Primary Head. We employ five disentangled heads to convert Gaussian embeddings (output from Gaussian Interaction Block) $\mathbf{E}$ into the corresponding five properties of 2D Gaussians $\mathcal{G}$. To be more specific, the $i$-th Gaussian embedding $\mathbf{E}_{i}$ is fed into five lightweight MLP blocks, each of which contains three linear layers and two ReLU activation functions, to output its opacity $\alpha_{i}$, color $c_{i}$, standard deviation $\sigma_{i}$, reference offset $o_{i}$, and correlation $\rho_{i}$. Note that the position is obtained by adding the reference position and predicted offset $\mu_{i}=p_{i}+o_{i}$. The five property heads are decoupled so that different heads can be adapted to predict specific properties. We apply the tanh function to ensure $\rho\in[-1,1]$, and use the sigmoid function for $\{\alpha,c,\sigma\}$ to ensure that their values are physically meaningful. No activation function is applied to $o$ since the positions of Gaussians in 2D images are expected to be free. Finally, the 2D Gaussians $\mathcal{G}$ are calculated from Eq. 1, Eq. 2 with the predicted five properties $\{\alpha,\mu,c,\sigma,\rho\}$.
@@
Scale
Computational Cost
and Performance
Backbone: EDSR-baseline